5, A mass m suspended from a string of elastic constant K oscillates with Amplitude A. If the amplitude of the oscillations is doubled, which of the following statement is correct?

A.     The frequency of the oscillation doubles

B.     The space traveled by the mass in a period doubles

C.     The oscillation period doubles

D.     The frequency of the oscillations is halved


1
Expert's answer
2021-09-09T17:34:28-0400

S - The space traveled by the mass in a period

A - Amplitude

s=4A, s1=4A, s2=4*2A=8A,

s2/s2=8A/4A=2,

it is showing that the space traveled by the mass in a period doubles.

Answer B is true.